X is 1/2 as old as Y. Z is 2 times as old as Y. Z is how many times as old as X?
Answer by oberobic(2304)

X = 1/2Y
Z = 2Y
.
redefine the equations in terms of Y = ??
.
X = 1/2Y
multiply by 2
2X = Y
Y = 2X
.
Z = 2Y
divide by 2
Z/2 = Y
Y = Z/2
.
Since they both = Y, we can put them = to one another.
therefore,
2X = Z/2
multiply by 2
4X = Z
.
so we can say Z is 4 times as old as X.
.
done